# https://krishnakanthnakka.github.io

### v1

@@ Abstract
Interpretability is critical in high-stakes domains such as medical imaging, where understanding model decisions is essential for clinical adoption. In this work, we introduce Sparse Autoencoder (SAE)-based interpretability to breast imaging by analyzing Mammo-CLIP, a vision–language foundation model pretrained on large-scale mammogram image–radiology report pairs. We train a patch-level Mammo-SAE on Mammo-CLIP visual features to identify and probe latent neurons associated with clinically relevant breast concepts such as mass and suspicious calcification. We show that top-activated class-level latent neurons often tend to align with ground-truth regions, and also uncover several confounding factors influencing the model’s decision-making process. Furthermore, we demonstrate that finetuning Mammo-CLIP leads to sharper concept separation in the latent space, improving interpretability and predictive performance. Our findings suggest that sparse latent representations offer a powerful lens into the internal behavior of breast foundation models.
@@ 2.2 Probing Mammo-SAE Latents to Identify Breast Concepts
For each class label $c\in\{0,1\}$, we compute the class-wise mean latent activation vector $\bar{\mathbf{z}}^{(c)}\in\mathbf{R}^{h}$ by averaging over all spatial locations and all training samples in that class:
@@ 3 Experiments
SAE Training. We utilize the Vision-SAEs library [19] to train a Sparse Autoencoder (SAE) on patch-level features extracted from the fine-tuned Mammo-CLIP model [5]. We focus specifically on the classifier trained for the suspicious calcification concept, using activations from the final layer of the EfficientNet-B5 backbone [9] of Mammo-CLIP [5]. Rather than training separate SAEs for each model which is expensive, we train a single SAE once and reuse it across all experiments. This shared SAE design not only reduces overhead but also ensures a consistent latent space, making it easier to compare representations across different models (see Section 3.3). The input feature dimension is $d=2048$, and we set the expansion factor to 8, resulting in a latent dimension of $h=16{,}384$. The SAE is trained for 200 epochs with a learning rate of $3\times 10^{-4}$, sparsity penalty $\lambda=3\times 10^{-5}$, and batch size of 4096.
@@ 3.1 Intervention on Class-level Latent Neurons
In the left panel of Figure 2(a), we show results for the top-$k$ activated intervention, where only the top-$k$ class-specific neurons are retained and all others are zeroed out, with $k$ varied from 0 up to the full latent dimensionality $h=16{,}384$. Remarkably, activating as few as 10 neurons is sufficient to nearly recover the model’s original AUC-ROC in multiple cases—demonstrating that a small subset of neurons captures most of the task-relevant signal.
@@ 3.2 Analyzing Top Activated Latent Neurons
Table 2: Mean Average Precision (mAP) for breast concept localization using top-$10$ class-level latent neuron activations of class $c=1$ across different breast cancer concepts and models.
@@ 4 Conclusion
5 Limitations
Our study has several limitations. First, we focus exclusively on the final layer of the Mammo-CLIP model, leaving the interpretability of earlier layers unexplored. Second, our evaluation is limited to only two breast cancer-related concepts—mass and suspicious calcification—and does not extend to other clinically relevant findings such as nipple retraction or skin thickening. Third, our analysis is confined to Mammo-CLIP; applying this interpretability framework to other vision-language models in medical imaging, such as MedCLIP [22] or GLoRIA [7], remains an important direction for future work. Finally, our visual probing is restricted to the top-10 class-level latent neurons per class, which may overlook other relevant or informative neurons.

### v2

@@ Abstract
Interpretability is critical in high-stakes domains such as medical imaging, where understanding model decisions is essential for clinical adoption. In this work, we introduce Sparse Autoencoder (SAE)-based interpretability to breast imaging by analyzing Mammo-CLIP, a vision–language foundation model pretrained on large-scale mammogram image–radiology report pairs. We train a patch-level Mammo-SAE on Mammo-CLIP visual features to identify and probe latent neurons associated with clinically relevant breast concepts such as mass and suspicious calcification. We show that top-activated class-level latent neurons often tend to align with ground-truth regions, and also uncover several confounding factors influencing the model’s decision-making process. Furthermore, we demonstrate that finetuning Mammo-CLIP leads to larger concept separation in the latent space, improving interpretability and predictive performance. Our findings suggest that sparse latent representations offer a powerful lens into the internal behavior of breast foundation models. The code will be released at https://krishnakanthnakka.github.io/MammoSAE/.
@@ 2.2 Probing Mammo-SAE Latents to Identify Breast Concepts
Given a dataset of $\{\mathbf{x},\mathbf{c}\}$ where each image $\mathbf{x}$ is annotated with a class label $c\in\{0,1\}$, where $c=1$ denotes the presence of the breast concept (eg., mass), we compute the class-wise mean latent activation vector $\bar{\mathbf{z}}^{(c)}\in\mathbf{R}^{h}$ by averaging over all spatial locations and all training samples in that class:
@@ 3 Experiments
SAE Training. We utilize the Vision-SAEs library [19] to train a Sparse Autoencoder (SAE) on patch-level features extracted from the fine-tuned Mammo-CLIP model [5]. We focus specifically on the classifier trained for the suspicious calcification concept, using activations from the final layer of the EfficientNet-B5 backbone [9] of Mammo-CLIP [5]. Rather than training separate SAEs for each model which is expensive, we train a single SAE once and reuse it across all experiments. This shared SAE design not only reduces overhead but also ensures a consistent latent space, making it easier to compare representations across different models (see Section 3.3). The input feature dimension is $d=2048$, and we set the expansion factor to 8, resulting in a latent dimension of h=16,384h=16{,}384. The SAE is trained for 200 epochs with a learning rate of $3\times 10^{-4}$, sparsity penalty $\lambda=3\times 10^{-5}$, and batch size of 4096.
@@ 3.1 Intervention on Class-level Latent Neurons
In the left panel of Figure 2(a), we show results for the top-$k$ activated intervention, where only the top-$k$ class-specific neurons are retained and all others are zeroed out, with $k$ varied from 0 up to the full latent dimensionality h=16,384h=16{,}384. Remarkably, activating as few as 10 neurons is sufficient to nearly recover the model’s original AUC-ROC in multiple cases—demonstrating that a small subset of neurons captures most of the task-relevant signal.
@@ 3.2 Analyzing Top Activated Latent Neurons
Table 2: Mean Average Precision (mAP) for breast concept localization using top-1010 class-level latent neuron activations of class $c=1$ across different breast cancer concepts and models.